All Exams Test series for 1 year @ ₹349 only
Question

P and Q are considering to apply for a job. The probability that P applies for the job is \(\frac{1}{4}\) , the probability that P applies for the job given that Q applies for the job is \(\frac{1}{2}\) , and the probability that Q applies for the job given that P applies for the job is \(\frac{1}{3}\) . Then the probability that P does not apply for the job given that Q does not apply for the job is

The correct answer is

\(\frac{4}{5}\)

Let A and B denote applications by P and Q. P(A∩B)=P(A)P(B|A)=(1/4)(1/3)=1/12. Hence P(B)=(1/12)/(1/2)=1/6. P(A′∩B′)=1−[1/4+1/6−1/12]=2/3 and P(B′)=5/6. Therefore P(A′|B′)=(2/3)/(5/6)=4/5.

Was this answer helpful?

Important Questions from Conditional Probability

  1. A box contains 5 coins: 4 regular coins and 1 fake coin. When a regular coin is tossed, the probability 𝑃(ℎ𝑒𝑎𝑑) = 0.5 and for a fake coin, 𝑃(ℎ𝑒𝑎𝑑) = 1. You pick a coin at random and toss it twice, and get two heads. The probability that the coin you have chosen is the fake coin is _______. (rounded off to two decimal places)

  2. A clinic specializes in testing for a disease D. The result of the test can be either positive or negative.

    A study revealed that if a person suffers from the disease D, the test result in that clinic comes out positive 80% of the time, and negative 20% of the time. If a person is not suffering from the disease D, the test comes out positive 10% of the time and negative 90% of the time. It is also known that among the general population, the disease D occurs in 30% of the individuals.

    If a person tests positive for D in that clinic, the probability that he/she actually suffers from the disease D is __________ . (Rounded off to two decimal places)
  3. Suppose an unbiased coin is tossed 6 times. Each coin toss is independent of all previous coin tosses. Let $E_1$ be the event that among the second, fourth, and sixth coin tosses, there are at least two heads. Let $E_2$ be the event that among the first, second, third, and fifth coin tosses, there are equal number of heads and tails.

    The conditional probability P($E_1$ | $E_2$) is equal to ____________. (rounded off to one decimal place)
  4. A box contains three apples and two oranges. Two fruits are removed randomly in succession. The probability that the first is an apple and the second is an orange is _____ .
  5. An experiment is performed by rolling an unbiased die once. Events A and B are defined as follows:
    Event A: The outcome of the rolled die is a prime number.
    Event B: The outcome of the rolled die is an odd number less than 4.
    The conditional probability of A given B, denoted by P(A|B), is ______ (rounded off to one decimal place).
Need Expert Advice?

Start Your Preparation with Prepp Mobile App

Download the app from Google Play & App Store
Download the app from Google Play & App Store
Prepp Mobile App